# VersiCode: Towards Version-controllable Code Generation

### v1

@@ B.1 Detailed Data Statistics of VersiCode
# Avg. Input Token 1,233 1,213 44 115 116 47 51 56
# Avg. Output Token 1 9 77 70 69 220 148 131
# Num. of Instances 13,532 13,530 1,618 49,346 49,346 32 21 82
B.2 Code Completion Datasets
As shown in Table 7, we compare the VersiCode-completion dataset with existing benchmarks. Note that, VersiCode stands out in annotated data size, marking it as the inaugural dataset tailored for version-specific generation.
@@ B.2 Code Completion Datasets
StaQC [48] StackOverflow Python, SQL 267,056 Function Programming Line-Level, Block-Level 2018 None
CoNaLa [49] StackOverflow Python, Java 2,879 Function Programming Line-Level, Block-Level 2018 Human
CT-maxmin [8] Existing Benchmark Multi(=6) 2,615 Cloze Test Token-Level 2020 None
HumanEval [6] Hand-Written Python 164 Function Programming Line-Level, Block-Level 2021 Human
MBPP [4] Hand-Written Python 974 Function Programming Block-Level 2021 Human
APPS [14] Programming Sites Python 10,000 Function Programming Line-Level, Block-Level 2021 None
CT-all [29] Existing Benchmark Multi(=6) 176,115 Cloze Test Token-Level 2021 None
CodeContests [25] Existing Benchmark, Codeforces Multi(=3) 13,610 Function Programming Block-Level 2022 None
HumanEval-FIM [9] Existing Benchmark Python 164 Function Programming Line-Level, Block-Level 2022 None
HumanEval+ [27] Existing Benchmark Python 164 Function Programming Line-Level, Block-Level 2023 Human
MBPP+ [27] Existing Benchmark Python 378 Function Programming Block-Level 2023 Human
DS-1000 [21] StackOverflow Python 1,000 Function Programming Line-Level, Block-Level 2023 Human
CoderEval [50] Github Python, Java 460 Function Programming Block-Level 2023 Human
@@
CodeApex [10] Programming Sites C++ 476 Function Programming Block-Level 2023 None
HumanEval-X [53] Existing Benchmark Multi(=5) 820 Function Programming Line-Level, Block-Level 2023 Human
VersiCode StackOverflow, Github Python, Java, C#, JavaScript 28,682 Cloze Test, Function Programming Token-Level, Line-Level, Block-Level 2024 Human, LLM
Appendix C Evaluation Details
C.1 Hyper-parameter
@@
Table 8: Hyper-parameters for completion and editing.
hyper-parameter code completion code editing
@@
C.2 Prompt Template
@@
Figure 9: Prompt template for version-aware code editing.
C.3 Data Sampling
For token and line-level tasks, we randomly sampled 2,000 instances for evaluation. For block-level tasks, we used the entire dataset due to its smaller size. In the time trend experiment, we sampled 200 data points per quarter or used all available data if fewer.
Appendix D Example Analysis
D.1 Error Analysis of GPT4-o
Despite GPT4-o achieving superior performance in general evaluation, it still encounters errors in 30% of instances. We provide several negative examples in Figure 10, Figure 11, Figure 12, and Figure 13.
@@ D.1 Error Analysis of GPT4-o
Figure 13: The fourth negative example of GPT-4o on token-level code completion.
Appendix E External Experiments
E.1 A Wider Comparison Study over LLMs
Table 9: Full evaluation results of Pass@1 on token-level code completion compared to related datasets and different data sources. The results for related datasets are collected from the online leaderboard of Evalplus [27].
Release Time Model HumanEval HumanEval+ MBPP MBPP+ VersiCode
Pass@1 Pass@1 Pass@1 Pass@1 Library Source Code Downstream Application StackOverflow Total
2023.06.14 WizardCoder-15B-V1.0 [32] 56.7 50.6 64.3 54.2 0.17 0 0.1 0.06
2023.06.14 WizardCoder-Python-7B-V1.0 [32] 50.6 45.1 58.5 49.5 6.62 0.17 5.45 2.66
2023.07.18 Llama-2-7B [42] 12.8 - 20.8 - 6.57 0.46 4.76 2.74
2023.07.18 Llama-2-13B-Chat [42] 18.3 - 30.6 - 3.71 0.06 3.41 1.51
2023.08.25 CodeLlama-7B-Instruct [37] 34.8 - 44.4 - 17.77 0.62 17.8 7.62
2023.08.25 CodeLlama-13B-Instruct [37] 42.7 - 49.4 - 28.45 2.47 32.05 13.5
2023.08.28 CodeLlama-7B-Python [37] 38.4 - 47.6 - 3.4 0.03 2.35 1.28
2023.10.29 DeepSeek-Coder-6.7B-Instruct [13] 74.4 71.3 74.9 65.6 3.83 0.15 4.34 1.71
2023.11.11 Mistral-7B-Instruct-V0.2 [17] 42.1 36 44.7 37 13.96 1.85 20.33 7.54
2024.01.25 DeepSeek-Coder-7B-Instruct-V1.5 [13] 75.6 71.3 75.2 62.2 26.7 4.51 44.77 15.71
2024.01.25 GPT-3.5-Turbo [34] 76.8 70.7 82.5 69.7 40.55 30.48 65.95 37.59
2024.02.27 StarCoder2-7B [28] 35.4 29.9 55.4 45.6 12.21 0.32 13.02 5.27
2024.02.27 StarCoder2-15B [28] 46.3 37.8 66.2 53.1 29.7 2.9 35.79 14.55
2024.04.09 CodeGemma-7B-Instruct [7] 60.4 51.8 70.4 56.9 31.8 0.76 31.29 13.36
2024.04.09 CodeGemma-7B [7] 44.5 41.5 65.1 52.4 29.61 1.12 34.01 13.28
2024.04.10 aiXCoder-7B [3] 54.9 - 66 - 17.51 1.09 26.3 8.83
2024.04.15 aiXCoder-7B-Base [3] 43.2 - 62.2 - 20.41 0.94 26.37 9.59
2024.04.15 CodeQwen1.5-7B [5] 51.8 45.7 73.5 60.8 11.61 0.12 7.58 4.33
2024.04.15 CodeQwen1.5-7B-Chat [5] 83.5 78.7 79.4 69 12.16 0.33 9.2 4.81
2024.04.18 Llama-3-8B [33] 35.5 29.3 61.4 51.6 17.18 0.24 20.69 7.57
2024.04.18 Llama-3-8B-Instruct [33] 61.6 56.7 70.1 59.3 20.79 3.67 34.08 12.23
2024.04.18 Llama-3-70B-Chat [33] 77.4 72 82.3 69 33.76 50.93 64.35 47.55
2024.05.13 GPT-4o [36] 85.4 81.7 85.7 73.3 58.37 72.98 87.21 70.44
E.2 Multi-language Analysis
Table 10: Multi-language performance on VersiCode
Model Python Java C# JavaScript
ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6
DeepSeek-Coder-7B-Instruct-V1.5 [13] 40.03 27.35 61.55 46.62 71.43 49.68 75.22 54.24
CodeLlama-13B-Instruct [37] 48.83 34.63 70.92 58.87 47.62 35.54 52.87 34.11
StarCoder2-15B [28] 39.71 27.36 38.63 27.43 33.33 28.63 60.67 39.33
CodeGemma-7B [7] 8.67 5.00 34.38 23.53 0 0 16.82 10.53
GPT-3.5-Turbo [34] 40.77 28.06 50.00 39.34 28.57 26.87 24.39 15.85
GPT-4o [36] 64.72 50.48 70.83 64.04 71.43 63.26 77.74 70.24
Llama-3-70B-Chat [33] 57.68 41.47 61.55 58.57 66.67 56.35 75.61 67.61
E.3 Block-level Generation without Grammar Verification
Figure 14: Python script for grammar verification.
As shown in Figure 14, we employ a straightforward yet effective function for basic grammar verification of generated Python code. Upon comparing Table 11 to Table 3, it becomes evident that the model tasked with editing, alongside reference code snippets from other versions, finds it easier to produce grammar-verified code.
Table 11: Block-level code completion and editing without grammar verification results.
Model Code Completion Code Editing (old to new) Code Editing (new to old)
Block-level Block-level Block-level
Score@6(ISM) Score@6(PM) Score@6(ISM) Score@6(PM) Score@6(ISM) Score@6(PM)
DeepSeek-Coder-7B-Instruct-V1.5 [13] 40.03 27.35 46.17 37.20 42.94 33.66
CodeLlama-13B-Instruct [37] 48.83 34.63 41.74 32.37 41.41 30.01
StarCoder2-15B [28] 39.71 27.36 40.94 30.73 44.46 31.88
CodeGemma-7B [7] 8.67 5.00 24.54 17.46 22.61 12.08
GPT-3.5-Turbo [34] 40.77 28.06 45.96 36.80 46.96 35.76
Llama-3-70B-Chat [33] 58.08 41.78 33.37 23.51 42.94 29.36
GPT-4o [36] 64.72 50.48 55.48 45.80 55.36 52.33

### v2

@@ Appendix B Data Statistics and Scope
# Avg. Input Token 2,087 2,075 55 191 195 57 63 67
# Avg. Output Token 2 16 128 131 128 349 255 167
# Num. of Instances 13,488 13,490 1,617 38,037 38,037 32 21 82
Scope: VersiCode supports version-specific code completion at the token, line, and block levels, enabling developers to navigate through version variations effortlessly. It also facilitates block-level version-aware code editing, empowering users to make precise modifications tailored to requirements of each version. The collected metadata also serves as a valuable resource for potential customized task modifications, supported domains are illustrated in Figure 8, aiding in fine-tuning workflows and enhancing model training for optimal performance.
Figure 8: A proportional chart based on the classification system of targeted audience and topics in third-party Python libraries on PyPI.
Appendix C Related Dataset
Code Completion Datasets. As shown in Table 5, we compare the VersiCode-completion dataset with existing benchmarks. VersiCode stands out in annotated data size, marking it as the inaugural dataset tailored for version-specific generation.
StaQC (Yao et al., 2018) StackOverflow Python, SQL 267,056 Function Programming Line-Level, Block-Level 2018 None
CoNaLa (Yin et al., 2018) StackOverflow Python, Java 2,879 Function Programming Line-Level, Block-Level 2018 Human
CT-maxmin (Feng et al., 2020) Existing Benchmark Multi(=6) 2,615 Cloze Test Token-Level 2020 None
HumanEval (Chen et al., 2021) Hand-Written Python 164 Function Programming Line-Level, Block-Level 2021 Human
MBPP (Austin et al., 2021) Hand-Written Python 974 Function Programming Block-Level 2021 Human
APPS (Hendrycks et al., 2021) Programming Sites Python 10,000 Function Programming Line-Level, Block-Level 2021 None
CT-all (Lu et al., 2021) Existing Benchmark Multi(=6) 176,115 Cloze Test Token-Level 2021 None
CodeContests (Li et al., 2022) Existing Benchmark, Codeforces Multi(=3) 13,610 Function Programming Block-Level 2022 None
HumanEval-FIM (Fried et al., 2023) Existing Benchmark Python 164 Function Programming Line-Level, Block-Level 2022 None
HumanEval+ (Liu et al., 2023) Existing Benchmark Python 164 Function Programming Line-Level, Block-Level 2023 LLM
MBPP+ (Liu et al., 2023) Existing Benchmark Python 378 Function Programming Block-Level 2023 LLM
DS-1000 (Lai et al., 2023) StackOverflow Python 1,000 Function Programming Line-Level, Block-Level 2023 Human
CoderEval (Yu et al., 2024) Github Python, Java 460 Function Programming Block-Level 2023 Human
CodeApex (Fu et al., 2023) Programming Sites C++ 476 Function Programming Block-Level 2023 None
HumanEval-X (Zheng et al., 2023) Existing Benchmark Multi(=5) 820 Function Programming Line-Level, Block-Level 2023 Human
BigCodeBench (Zhuo et al., 2024) Existing Benchmark Python 1,140 Function Programming Block-Level 2024 Human, LLM
VersiCode StackOverflow, Github Python, Java, C#, JavaScript 28,595 Cloze Test, Function Programming Token-Level, Line-Level, Block-Level 2024 Human, LLM
@@ Appendix C Related Dataset
Code Migration Datasets. As shown in Table 6, we compare the VersiCode-migration dataset with existing benchmarks. VersiCode stands out in annotated data size, marking it the inaugural dataset tailored for version-specific migration.
Defects4J (Just et al., 2014) Open Source Programs Java 357 Debug Block-Level 2014 None
QuixBugs (Lin et al., 2017) Quixey Challenges Python, Java 40 Debug Line-Level 2017 Human
CoST (Zhu et al., 2022b) GeeksForGeeks Multi(=7) 132,046 Code Translation Line-Level, Block-Level 2022 None
XLCoST (Zhu et al., 2022a) GeeksForGeeks Multi(=8) 1,083,000 Code Translation Line-Level, Block-Level 2022 None
InstructCoder (Hu et al., 2023) Github Python 114,000 Code Refinement Block-Level 2023 LLM
@@
MultilingualTrans (Yan et al., 2023) Programming Sites Multi(=8) 30,419 Code Translation Block-Level 2023 None
NicheTrans (Yan et al., 2023) Programming Sites Multi(¿8) 236,468 Code Translation Block-Level 2023 None
LLMTrans (Yan et al., 2023) Hand-Written Multi(=8) 350 Code Translation Block-Level 2023 Human
Avatar (Ahmad et al., 2023) Programming Sites Python, Java 62,520 Code Translation Block-Level 2023 None
G-TransEval (Jiao et al., 2023) Existing benchmark, GeeksForGeeks Multi(=5) 400 Code Translation Token-Level, Block-Level 2023 Human
EvalGPTFix (Zhang et al., 2023) AtCoder Java 151 Debug Block-Level 2023 Human
DebugBench (Tian et al., 2024) LeetCode Multi(=3) 4,253 Debug Block-Level 2024 LLM
VersiCode Github Python 76,074 Version Adaptation Block-Level 2024 LLM
Appendix D Additional Experiments and Details
D.1 Extensive Comparative Study on Large Language Models
Release Time Model HumanEval HumanEval+ MBPP MBPP+ VersiCode
EM@1 EM@1 EM@1 EM@1 Library Source Code Downstream Application StackOverflow Total
2023.06.14 WizardCoder-15B-V1.0 (Luo et al., 2024c) 56.7 50.6 64.3 54.2 0.17 0 0.1 0.06
2023.06.14 WizardCoder-Python-7B-V1.0 (Luo et al., 2024c) 50.6 45.1 58.5 49.5 6.62 0.17 5.45 2.66
2023.07.18 Llama-2-7B (Touvron et al., 2023) 12.8 - 20.8 - 6.57 0.46 4.76 2.74
2023.07.18 Llama-2-13B-Chat (Touvron et al., 2023) 18.3 - 30.6 - 3.71 0.06 3.41 1.51
2023.08.25 CodeLlama-7B-Instruct (Rozière et al., 2023) 34.8 - 44.4 - 17.77 0.62 17.8 7.62
2023.08.25 CodeLlama-13B-Instruct (Rozière et al., 2023) 42.7 - 49.4 - 28.45 2.47 32.05 13.5
2023.08.28 CodeLlama-7B-Python (Rozière et al., 2023) 38.4 - 47.6 - 3.4 0.03 2.35 1.28
2023.10.29 DeepSeek-Coder-6.7B-Instruct (Guo et al., 2024) 74.4 71.3 74.9 65.6 3.83 0.15 4.34 1.71
2023.11.11 Mistral-7B-Instruct-V0.2 (Jiang et al., 2023) 42.1 36 44.7 37 13.96 1.85 20.33 7.54
2024.01.25 DeepSeek-Coder-7B-Instruct-V1.5 (Guo et al., 2024) 75.6 71.3 75.2 62.2 26.7 4.51 44.77 15.71
2024.01.25 GPT-3.5-Turbo (OpenAI, 2023b) 76.8 70.7 82.5 69.7 40.55 30.48 65.95 37.59
2024.02.27 StarCoder2-7B (Lozhkov et al., 2024) 35.4 29.9 55.4 45.6 12.21 0.32 13.02 5.27
2024.02.27 StarCoder2-15B (Lozhkov et al., 2024) 46.3 37.8 66.2 53.1 29.7 2.9 35.79 14.55
2024.04.09 CodeGemma-7B-Instruct (CodeGemma Team et al., 2024) 60.4 51.8 70.4 56.9 31.8 0.76 31.29 13.36
2024.04.09 CodeGemma-7B (CodeGemma Team et al., 2024) 44.5 41.5 65.1 52.4 29.61 1.12 34.01 13.28
2024.04.10 aiXCoder-7B (aiXcoder team, 2024) 54.9 - 66 - 17.51 1.09 26.3 8.83
2024.04.15 aiXCoder-7B-Base (aiXcoder team, 2024) 43.2 - 62.2 - 20.41 0.94 26.37 9.59
2024.04.15 CodeQwen1.5-7B (Bai et al., 2023) 51.8 45.7 73.5 60.8 11.61 0.12 7.58 4.33
2024.04.15 CodeQwen1.5-7B-Chat (Bai et al., 2023) 83.5 78.7 79.4 69 12.16 0.33 9.2 4.81
2024.04.18 Llama-3-8B (Meta LlaMa team, 2024) 35.5 29.3 61.4 51.6 17.18 0.24 20.69 7.57
2024.04.18 Llama-3-8B-Instruct (Meta LlaMa team, 2024) 61.6 56.7 70.1 59.3 20.79 3.67 34.08 12.23
2024.04.18 Llama-3-70B-Chat (Meta LlaMa team, 2024) 77.4 72 82.3 69 33.76 50.93 64.35 47.55
2024.05.13 GPT-4o (OpenAI, 2024) 85.4 81.7 85.7 73.3 58.37 72.98 87.21 70.44
Table 7: Full evaluation results of EM@1 on token-level code completion compared to related datasets and different data sources. The results for related datasets are collected from the online leaderboard of Evalplus (Liu et al., 2023).
D.2 Multi-language Analysis
Model Python Java C# JavaScript
ISM@1 PM@1 ISM@1 PM@1 ISM@1 PM@1 ISM@1 PM@1
DeepSeek-Coder-7B-Instruct-V1.5 (Guo et al., 2024) 40.03 27.35 61.55 46.62 71.43 49.68 75.22 54.24
CodeLlama-13B-Instruct (Rozière et al., 2023) 48.83 34.63 70.92 58.87 47.62 35.54 52.87 34.11
StarCoder2-15B (Lozhkov et al., 2024) 39.71 27.36 38.63 27.43 33.33 28.63 60.67 39.33
CodeGemma-7B (CodeGemma Team et al., 2024) 8.67 5.00 34.38 23.53 0 0 16.82 10.53
GPT-3.5-Turbo (OpenAI, 2023b) 40.77 28.06 50.00 39.34 28.57 26.87 24.39 15.85
GPT-4o (OpenAI, 2024) 64.72 50.48 70.83 64.04 71.43 63.26 77.74 70.24
Llama-3-70B-Chat (Meta LlaMa team, 2024) 57.68 41.47 61.55 58.57 66.67 56.35 75.61 67.61
Table 8: Multi-language performance on VersiCode
D.3 Block-level Generation without Grammar Verification
We use Python’s built-in function “compile()” to compile the generated code snippets in order to check whether they are syntactically correct. Upon comparing “w/o grammar verification” and “w grammar verification” in Table 9, it becomes evident that the model tasked with editing, alongside reference code snippets from other versions, finds it easier to produce grammar-verified code.
Table 9: Results of block-level code completion and migration with or without grammar verification.
Appendix E Metric Design of Critical Diff Check
E.1 Introduction of Critical Diff Check
Critical Diff Check (CDC) focuses on the changes in the code rather than the overall similarity of the entire code segment. CDC has five rules as follows:
  • Rule 1: Check whether the generated code contains the core token.
  • Rule 2: Check whether the generated code is valid.
  • Rule 3: Check if the number of arguments in the function using the core token is consistent.
  • Rule 4: If the reference code uses a with statement, checks whether the generated code also uses a with statement.
  • Rule 5: If the reference code uses keyword argument assignment, checks whether the generated code uses the same keyword argument assignment.
The failure frequency and examples for each rule are shown in Table 10.
Table 10: Each rule of the CDC, along with the frequency, occurrence rate, and examples of mismatches for each rule. ‘a’ represents the core token, ‘c’ represents the code generated by the model, ‘c’ represents the reference code, ‘f’ represents the function of the specified token, and ‘params’ refers to the function’s parameter list. ‘Kc(f)’ and ‘Kc(f)’ represent the keyword parameter lists of the reference code and the model-generated code, respectively, and ‘p’ represents the parameter assigned using keyword arguments. In detail, Rule 1 checks whether the generated code contains the core token; Rule 2 checks whether the generated code is valid; Rule 3 checks if the number of arguments in the function using the core token is consistent; Rule 4, if the reference code uses a with statement, checks whether the generated code also uses a with statement; Rule 5, if the reference code uses keyword argument assignment, checks whether the generated code uses the same keyword argument assignment.
E.2 Ablation Study of Critical Diff Check
We conducted ablation experiments on the five CDC rules and calculated the Pearson correlation coefficient with the Pass@1 metric for each, to demonstrate the reliability of CDC. The specific experimental data is shown in Table 11.
Table 11: Ablation study of Critical Diff Check per rule. The configuration labeled as “CDC w/o Rule i”, where $i\in\{1,2,3,4,5\}$ means that when calculating the CDC score, Rule i is excluded, and only the other four rules are considered. The Pearson correlation coefficient calculates the correlation the metric’s results obtained in each configuration against Pass@1.
Appendix F Running Example of Executable Test
As shown in Figure 9, this is an example of a task function used for code generation, where the task function is processed in various granular forms of code completion. The “core token” is only provided for visualization, which is unseen for models. “library version” is optional, identified as “w/ or w/o version”, and “import” statements are also optional, identified as “w/ or w/o import” in Table 1. As shown in Figure 10, these are the test cases for the task function illustrated in Figure 9. The test cases were developed by experts through interactions with GPT-4 and include four types of tests.
Figure 9: The ground truth for block-level code generation, used for Section 4. Note that, “core token” is only provided for visualization, which is unseen for models. “library version” is optional, identified as “w/ or w/o version”, and “import” statements are also optional, identified as “w/ or w/o import” in Table 1.
Figure 10: The test cases associated with generated code for dynamic code analysis, used for Section 4.
Appendix G Evaluation Details
G.1 Hyper-parameter
@@
hyper-parameter code completion code migration
@@
Table 12: Hyper-parameters for completion and migration.
G.2 Prompt Template
@@
Figure 14: Prompt template for version-aware code migration.
G.3 Data Sampling
For token-level completion tasks(Figure 3), we randomly sampled 2,000 instances for evaluation. We used the entire executable dataset for line- and block-level completion tasks due to its smaller size (Figure 6, Table 1). In the time trend experiment (Figure 4), we sampled 200 data points per quarter or used all available data if fewer. And in the code migration task (Table 2), we randomly sampled 2,000 instances for evaluation.
Appendix H Error Analysis
H.1 Error Analysis of GPT4-o
Despite GPT4-o achieving superior performance in general evaluation, it still encounters errors in 30% of instances. We provide several negative examples in Figure 15, Figure 16, and Figure 17.